one leg of a right triangle is 7 units longer than the length of the other leg if the hypotenuse is 13 then find the other two legs decide what to make x write and equation would it be 13=7+2(x) with x=3?

Answer by rajagopalan(174)

Let one leg = x
other leg=x+7
then x^2+(x+7)^2=13^2
x^2+x^2+14x+49=169
2x^2+14x+49=169
2x^2+14x+49-169=0
2x^2+14x-120=0
x^2+7x-60=0 divide both sides by 2
x^2+12x-5x-60=0
x(x+12)-5(x+12)=0
(x-5)(x+12)=0
giving x=5 or -12. ignore negative value
we get x=5
other side=5+7=12
okay